# From weak to strong-coupling superconductivity tuned by substrate in TiN films

### v1

@@ V Conclusions
Superconducting properties are investigated in TiN films grown on two different substrates. The superconducting critical temperature, normal-state resistivity and upper critical field all reveal a systematically evolution with the disorder level. Strikingly, the system undergoes a transition from weak coupling ($\Delta_{0}/k_{B}T_{c}$ = 1.60 $\sim$ 1.75) to strong coupling ($\Delta_{0}/k_{B}T_{c}$ = 3.05), accompanied by a change in film orientation from (111) to (100), as the substrate changes from sapphire to Si. Further analysis indicates that this change in superconducting coupling strength is closely related to the difference in Coulomb screening effect caused by disorder and the change in density of states at the Fermi level.
VI Data availability statement
@@
VII ACKNOWLEDGMENTS
This work is supported by the Shanghai Technology Innovation Action Plan Integrated Circuit Technology Support Program (Grant No. 22DZ1100200), and the National Key Research and Development Program of China (Grant No. 2023YFB4404904). The authors would like to thank all the staff at the Superconducting Electronics Facility (SELF) for their assistance.

### v2

@@ 5. Conclusions
Superconducting properties are investigated in TiN films grown on three different substrates. The SC critical temperature, normal-state resistivity and upper critical field all reveal a systematically evolution with the disorder level. Strikingly, the system undergoes a transition from weak coupling ($\Delta_{0}/k_{B}T_{c}$ = 1.60 $\sim$ 1.75) to strong coupling ($\Delta_{0}/k_{B}T_{c}$ = 3.05), accompanied by a change of the substrate changes from sapphire and MgO to Si. Further analysis indicates that this change in superconducting coupling strength could not be fully explained by the disorder level or film orientations. The surface morphology, especially the grain size, of the thin films may be an important factor influencing the superconductivity.
Data availability statement
@@
ACKNOWLEDGMENTS
This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0670000), the National Key Research and Development Program of China (Grant No. 2023YFB4404904), the Key-Area Research and Development Program of Guangdong Province, China (No. 2020B0303030002), and the Autonomous Deployment Project of State Key Laboratory of Materials for Integrated Circuits (No. SKLJC-Z2024-B04). The authors would like to thank all the staff at the Superconducting Electronics Facility (SELF) for their assistance.